OpenStudy (anonymous):

what is the average value of 78,83,65

OpenStudy (anonymous):

To find the average, sum all the terms and divide that by the number of terms. Average of 1 and 3: (1+3)/2 = 4/2 = 2.

OpenStudy (anonymous):

oohhh thanks